OpenStudy (anonymous):

Find the standard form of the equation of the hyperbola with the given characteristics. foci= (+/-3,0) asymptotes= +/-2x

OpenStudy (anonymous):

asymptotes: y=\[\pm2x\]

OpenStudy (mertsj):

a^2+b^2=c^2 for a hyperbola. c=3 b/a=2 b^2=4a^2 a^2+4a^2=9 a^2=9/5 b^2=36/5 Equation: \[\frac{x ^{2}}{\frac{9}{5}}-\frac{y ^{2}}{\frac{36}9}\]

OpenStudy (mertsj):

=1 of course